THe equation of line k is 2x-5y=-3. Write the equation of the line that passes through the point (-1,4) and is perpendicular to line k.
Answer by scott8148(6628)

perpendicular lines have slopes that are negative reciprocals of each other

from the standpoint of algebraic mechanics , the easiest technique is to swap the x and y coefficients and change one of the signs

a line perpendicular to 2x-5y=-3 is ___ 5x+2y=?

to find the ? , just plug in the coordinates of a point included in the line

for (-1,4) ___ 5(-1)+2(4)=? ___ -5+8=? ___ 3=?

so the new equation is ___ 5x+2y=3